### v3

# Mirage of Mastery: Memorization Tricks LLMs into Artificially Inflated Self-Knowledge

Abstract
When artificial intelligence mistakes memorization for intelligence, it creates a dangerous mirage of reasoning. Existing studies treat memorization and self-knowledge deficits in large language models (LLMs) as separate issues and do not recognize an intertwining link that degrades the safety and trustworthiness of LLM responses. In our study, we utilize a novel framework to ascertain if LLMs genuinely learn reasoning patterns from training data or merely memorize them to assume competence across problems of similar complexity focused on STEM domains. Our analysis shows a noteworthy problem in generalization: LLMs draw confidence from memorized solutions to infer a higher self-knowledge about their reasoning ability, which manifests as an over 45% inconsistency in feasibility assessments when faced with validated, logically coherent task perturbations. This effect is most pronounced in science and medicine domains, which tend to have maximal standardized jargon and problems, further confirming our approach. Such overconfidence poses critical AI safety risks in trust-sensitive domains, like healthcare, law, and scientific research, where unreliable responses can have severe consequences. Significant wavering within the self-knowledge of LLMs also shows flaws in current architectures and training patterns, highlighting the need for techniques that ensure a balanced, consistent stance on models’ perceptions of their own knowledge. Our code and results are available publicly.  [^1] https://github.com/Sahil-R-Kale/mirage_of_mastery
1 Introduction
For true reliability and trustworthiness, AI tools must consistently and accurately recognize the boundary of their own capabilities, referred to as self-knowledge [yin2023largelanguagemodelsknow]. In the case of large language models (LLMs), it refers to their introspective ability to determine which tasks and prompts are feasible precisely and with certainty. [ni2024llmsneedretrievalaugmentation, kale2025linedutyevaluatingllm]. Previous methodologies studying LLM self-knowledge are rife [wang-etal-2023-self-knowledge, wen2024perceptionknowledgeboundarylarge, ren2024investigatingfactualknowledgeboundary], and have shown that LLMs lack a grounded understanding of their own knowledge boundaries. However, only rarely have other challenges like memorization [slonski2024detectingmemorizationlargelanguage] and adversarial helpfulness [ajwani2024llmgeneratedblackboxexplanationsadversarially] been attributed as underlying causes for inaccuracies and overconfidence in self-knowledge, a gap needing timely redress.
Figure 1: Broad idea of memorization-driven skew in self-knowledge: inconsistent feasibility assessments with minor task perturbations (underlined). The task on the left is generated by the LLM itself with a confident claim of answerability, while slight perturbations on the right are then deemed infeasible.
Recent research highlights two key challenges in large language models: memorisation, the tendency to recall and reproduce training data without true understanding [satvaty2025undesirablememorizationlargelanguage, prashanth2024recitereconstructrecollectmemorization], and, as previously described, a lack of self-knowledge, or the inability to accurately demarcate the limits of their own capabilities [tan2024iunderstandicreate]. As data used for LLM training scales to include more problems and solutions, especially in STEM fields, we believe models may conflate the ability of ‘knowing’ solutions with genuine reasoning power to solve such problems. This memorization-driven behaviour raises a critical downstream concern regarding generalization: if LLMs mistake recall for reasoning, they may falsely perceive their self-knowledge to be deeper than it truly is, as shown in Figure 1. Such overconfidence, drawn from memorized problem data and patterns, can lead to unreliable responses, harming AI trustworthiness and safety.
Our inspiration can be explained with a simple human parallel: a medical student who successfully diagnoses and treats one patient will not assume they can automatically solve any other case of similar difficulty without understanding the underlying medical principles. In contrast, we find that LLMs tend to behave as if they do, drawing undue confidence from memorized examples. We propose a novel, universally applicable framework to determine whether LLMs genuinely learn reasoning patterns from training data or merely memorize them, thereby erroneously assuming competence across problems of similar complexity.
In our methodology, we give LLMs the freedom to provide a problem in STEM fields they are confident of solving and evaluate their consistency in feasibility analyses on perturbed problems of similar complexity, requiring the same reasoning patterns. If LLMs rely on memorization for self-knowledge, their accuracy and consistency in answering and determining the feasibility of tasks should falter when faced with minor logical perturbations [xie2025memorizationlargelanguagemodels]. We thus develop a method to identify such inconsistencies in feasibility assessments of logically coherent task variations, revealing if and when LLMs rely on memorization to artificially inflate and generalize their self-knowledge.
Our analysis shows that LLMs indeed draw confidence from memorized solutions to infer a higher self-knowledge about their reasoning ability, showing a significant gap in generalization. Models tend to mistake recall for reasoning, and self-knowledge assessments are largely influenced by prior exposure to problems rather than a genuine understanding of problem-solving ability. Moreover, even powerful models display vast inconsistencies in feasibility judgments for structurally similar problems. Such misplaced trust in AI reasoning in high-stakes fields like science, law, and medicine can lead to critical consequences. Our key contributions can be summarized as follows:
  1. 1.
    We identify a key link between two known problems of language models that act together to degrade the reliability of LLM responses and present an effective method to analyse the same
  2. 2.
    To quantify the interplay between self-knowledge and memorization, we provide a universally applicable task perturbation pipeline and metrics analysing these results
  3. 3.
    We show how overconfidence and a lack of true reasoning awareness in powerful LLMs stems from memorized solutions and data, and observe a significant lack of consistency in self-knowledge judgements
Figure 2: Methodology to analyse how LLM memorization inflates self-perception using minor task perturbations
2 Experimentation Methodology
We present a novel experimentation approach to analyse if LLMs’ tendency to memorise problems and corresponding solutions builds an inflated perception of their own knowledge and capabilities. Our methodology builds on the self-knowledge evaluation technique given by [kale2025linedutyevaluatingllm] and the principle from research by [xie2025memorizationlargelanguagemodels] that memorization is characterized by high accuracy on familiar problems but a sharp decline when faced with slight problem variations.
To effectively measure the interplay between memorization and self-knowledge, we devise a structured process to (1) generate tasks that LLMs consider feasible, either due to memorization of solutions or confidence in reasoning ability and (2) evaluate shifts in self-knowledge assessments when faced with perturbed versions of the same tasks that preserve problem difficulty. In short, we develop a method to identify inconsistencies in feasibility assessments of self-validated, logically coherent task variations, revealing when LLMs rely on memorization to exaggerate their self-knowledge. Towards this goal, we design a dynamic experimentation approach presented in Figure 2.
Table 1: MIRAGE scores for all models distributed across STEM domains
Model Total $\downarrow$ Science Technology Engineering Medicine
GPT-4o 0.79 0.85 0.82 0.66 0.84
DeepSeek-V3 0.79 0.83 0.88 0.64 0.79
Mistral Large 24.11 0.58 0.96 0.42 0.55 0.40
Claude 3.7 Sonnet 0.46 0.73 0.31 0.37 0.42
Table 2: SKEW scores for all models distributed across STEM domains
Model Total $\downarrow$ Science Technology Engineering Medicine
GPT-4o 0.51 0.50 0.53 0.50 0.45
DeepSeek-V3 0.51 0.53 0.55 0.51 0.46
Mistral Large 24.11 0.35 0.34 0.51 0.40 0.42
Claude 3.7 Sonnet 0.32 0.42 0.48 0.40 0.36
2.1 Task Generation
In the initial stage, we use a structured prompt to generate tasks with task instructions and associated data, if any, with the constraint that the LLM finds them completely feasible. During the generation phase, we allow LLMs the flexibility to set their own feasibility boundaries and use the QAP [yugeswardeenoo-etal-2024-question] approach while devising prompts to ensure introspection in self-knowledge before providing responses. While we restrict the task domain to STEM fields, no other constraints are imposed in the prompt for task generation (refer to the prompt in Figure 6 in Appendix Appendix). Since LLMs are prone to inconsistencies in self-knowledge assessments [kale2025linedutyevaluatingllm], we incorporate a separate prompt-based validation step (refer to the prompt in Figure 7 in Appendix Appendix) within the task generation phase to ensure that models are highly confident in their ability to answer the problems they have generated. Also, we conduct spot checks using STEM field experts to ensure task feasibility. We feed such validated feasible tasks to the task perturbation pipeline ahead.
2.2 Task Perturbation Pipeline
In order to check whether LLMs’ self-knowledge about reasoning capabilities is indeed memorization-driven, we design a pipeline to add minor perturbations to the feasible tasks generated in the previous step while maintaining complete logical cohesion, similar computational complexity and domain relevance. Our pipeline consists of 3 main modules implemented sequentially. We iteratively refined the pipeline through manual spot checks on 10% of the samples, ensuring that the perturbed tasks were logically sound and maintained feasibility.
Ontology replacement consists of substituting key domain-specific terms in the original task instructions and data with equivalent terminology from the same STEM domain. We also modify relationships between entities in the task while preserving logical coherence in associated data. This step is crucial to minimize terminology-driven memorization while preserving the task’s core formulation and structure. So as to also incorporate basic language level perturbation, we implement this module by setting up a Gemini 1.5 Flash model [geminiteam2024gemini15unlockingmultimodal] with specific replacement instructions in the prompt shown in Figure 8 in Appendix Appendix.
The instruction translation module involves changing the language of the English task instructions to any of German, Spanish or French using the Google Translate API. We select these languages since LLMs themselves indicate they have the most training resources for them and demonstrate the highest confidence in their understanding. By only translating instructions and not data, we avoid inconsistencies in domain-specific knowledge, which is often best understood by models in English.
The data perturbation module uses a simple rule-based approach to introduce an approximately 15 percent variation in all numerical values present in the task data. Since LLMs are prone to memorization of numerical patterns and data [bordt2024elephantsforgetmemorizationlearning], this step is crucial to dissociate tasks from such patterns. We also reorder all unordered data elements including lists and arrays in the task data to minimize pattern matching while answering and classifying feasibility of tasks.
For all perturbed tasks, to ensure feasibility, we take the help of STEM field experts to manually verify feasibility and remove samples where difficulty level is increased due to language vocabulary limitations, or data-related inconsistencies.
2.3 Task Classification
In the final stage, we feed all perturbed tasks generated through the pipeline to the LLM to attempt (refer to the prompt in Figure 9 in Appendix Appendix). For each task, the LLM is prompted to either generate a conclusive answer (and thus classify the task as feasible) or mark it as infeasible if it finds the task to lie beyond its self-identified reasoning capabilities. Since all tasks are validated to be feasible and are perturbed in a way that maintains domain relevance and complete logical cohesion, an inconsistency in feasibility assessments strongly implies that the LLMs relied on memorization of the data or solution steps of the original task to draw confidence for an inflated sense of self-knowledge and reasoning capacity in that domain.
3 Evaluation
3.1 Formulation and Metrics
To formulate our approach mathematically, we can describe it as follows. First, we prompt an LLM to generate a task $\tau{}$, where $\tau{}$ is explicitly deemed feasible by the model ($F(\tau)=1$ for feasible tasks) and confirmed via another self-validation. We repeat this process $m$ times to get $m$ distinct tasks ($\{\tau_{1},\tau_{2},\dots,\tau_{m}\}$). Each generated task $\tau_{i}$ is then perturbed $n$ times via the perturbation pipeline that modifies surface-level features while ensuring logical cohesion and equivalent problem complexity. This produces the perturbed task set: $\{P_{1}(\tau_{i}),P_{2}(\tau_{i}),\dots,P_{n}(\tau_{i})\}$.
The perturbed task set is then re-evaluated by the LLM, where each task is either classified as feasible or infeasible. An inconsistency in classification, i.e., $F(\tau_{i})\neq F(P_{j}(\tau_{i}))$, indicates that the model relied on memorization rather than genuine reasoning ability to assess the feasibility of the original task, leading to overconfidence in self-knowledge.
Owing to the novelty of exploring such a relation between known LLM problems, and a lack of suitable existing benchmarks, we devise two new metrics based on simple averages to quantify our experimental results.
To quantify the proportion of times how often an LLM changes its feasibility stance about perturbed tasks originally generated by itself as feasible, we introduce a new metric called MIRAGE (Memorization Induced Reasoning Assumption & GEneralization). It represents the mean infeasibility rate across perturbed tasks, averaged over all sets. A high MIRAGE score shows a high proportion of flipping judgements, implying that models considered the original task to be feasible majorly because of memorized patterns.
$\text{MIRAGE}=\frac{1}{m}\sum_{i=1}^{m}\frac{\sum_{j=1}^{n}\mathds{1}(F(P_{j}(\tau_{i}))=0)}{n}$ (1)
Similarly, our research findings can also be used to methodically quantify and analyse consistency in self-knowledge of LLMs across various domains. For this purpose, we propose a metric called SKEW (Self-KnowlEdge Wavering). SKEW quantifies the inconsistency in feasibility assessment for very similar problems with minor perturbations, such that a higher score implies lower agreement, indicating poor self-knowledge about feasibility boundaries. Unlike MIRAGE, here we include the original task along with its $n$ perturbations, resulting in $t=n+1$ tasks per set $S_{i}$, for $i=1,\dots,m$.
$\text{SKEW}=\frac{1}{m}\sum_{i=1}^{m}\frac{\sum_{\begin{subarray}{c}\tau_{a},\tau_{b}\in S_{i}\ \tau_{b}>\tau_{a}\end{subarray}}\mathds{1}(F(\tau_{a})\neq F(\tau_{b}))}{\binom{t}{2}}$ (2)
3.2 Setup
For comprehensive analysis, we experiment with a wide range of high-performance models. Since our methodology is applicable across both closed and open-source models, we choose 2 closed-source models, GPT-4o [openai2025gpt4o] and Claude 3.7 Sonnet [anthropic2025claude3], and 2 open-source models, Mistral Large 24.11 [mistral2025large2407] and DeepSeek v3 [deepseekai2024deepseekv3], in our evaluation. For all models, detailed model parameters are provided in Section .2 in Appendix Appendix. For each STEM domain and model, we generate 34 original tasks and perturb each task 3 times, leading to a dataset of 102 perturbed tasks per domain and 408 across all domains. MIRAGE scores across LLMs for all domains are presented in Table 1, while SKEW values analysing self-knowledge inconsistencies are given in Table 2.
Figure 3: MIRAGE scores for LLMs across STEM domains, with the overall average baseline represented in red
Figure 4: SKEW scores for LLMs across STEM domains, with the overall average baseline represented in red
Figure 5: Results showing LLM performance metrics measuring memorization-driven self-knowledge inflation
4 Quantifying Memorization-inflated Self-Knowledge
Self-knowledge inflation due to memorization is striking. The consistently high MIRAGE scores for all models, as shown in Table 1 and visualized in Figure 3, point to an alarming fundamental flaw: models are not just mistaking recall for reasoning but also inflating their self-knowledge perception using this memorization-driven confidence. Even the best high-performance models like GPT-4o and Mistral Large 24.11 change their feasibility stance about slightly perturbed tasks over 45% of time, meaning that LLMs are systematically overestimating their ability based on memorized solutions and data rather than true reasoning awareness. A likely cause of this widespread memorization-induced overconfidence could be the prevalence of STEM benchmarks and evaluations in training data, which may act more towards reinforcing memorization patterns rather than fostering robust reasoning in LLMs. This finding highlights a significant trust issue and a critical vulnerability in AI reliability and generalization.
Science and medicine are LLM memorization hotspots. Almost all models show the highest MIRAGE scores for the science and medicine domains, with Mistral’s extreme score of 0.96 exposing corpus-specific overfitting. Since fields like science and medicine tend to have the most frequency of standardized jargon and problem formats, models may draw overconfidence in reasoning abilities from such patterns. It is likely that repeated textbook-style problems in training data exacerbate this problem. Since applications in these fields are generally the most critical, future training data should diversify science and medicine-related texts to ensure a stable outlook towards reasoning capacity.
Architectural choice alone is seemingly insufficient to regulate memorisation-inflated self-knowledge risk consistently across domains. The general inconsistency in MIRAGE values across both domains and models suggests that a specific model or training architecture is not sufficient to manage risks of artificially increased self-knowledge across STEM domains. This challenges the notion that architectural improvements or scaling alone improve generalization and consistency. There is a need for other advancements and algorithms that mitigate overconfidence during training and ensure a stable outlook towards self-knowledge of reasoning capabilities across domains.
5 Impact of Self-Knowledge Inconsistencies
LLMs lack the capacity to establish generalizable feasibility boundaries. High SKEW values seen in Table 2 and Figure 4, particularly for GPT-4o and DeepSeek-V3, indicate that even minor perturbations disrupt feasibility judgments, highlighting that models lack a generalized, consistent stance of their own reasoning ability, especially in STEM domains. While Mistral and Claude exhibit comparatively greater stability, peaks nearing 0.5 SKEW suggest that current architectures and training patterns do not suffice in establishing a consistent stance on models’ perceptions of their own knowledge.
Unstable self-knowledge reflects memorization dependence The high MIRAGE values across all LLMs show that memorization leads to an overestimation in self-knowledge about STEM reasoning. Models do not possess a stable internal metric for their own reasoning capability and instead rely on memorized patterns and data that collapse equally under perturbation.
LLMs are not trustworthy enough for critical real-world applications. Our findings highlight fundamental challenges and potential advancements in deploying LLMs for high-stakes domains like healthcare, law, and research, where memorized or unreliable reasoning can cause critical real-world consequences. We demonstrate how different LLMs exhibit domain-specific weaknesses across distinct STEM fields and recommend adaptive LLM routing strategies [ong2024routellmlearningroutellms] to be wary of these issues during the selection of models for important tasks. Given that most models exhibit artificially inflated self-knowledge in science and medicine due to memorization, we emphasize the need for cautious deployment in these fields. Implementing safeguards like confidence thresholds and source markers can help flag uncertain responses, ensuring users are aware of potential inaccuracies before using AI-generated outputs elsewhere. Due to their inflated self-perception, models are prone to generating responses even when lacking sufficient knowledge, rather than abstaining. Hence, human-in-the-loop fallback strategies are still important in LLM-powered applications for maximum trustworthiness. We urge future research to prioritize methods that enhance self-knowledge robustness, reduce over-reliance on memorization, and establish mechanisms for models to recognize and convey lack in their own knowledge.
6 Related Work
Memorization in LLMs: Memorization, be it verbatim or contextual, has garnered significant attention due to its implications for privacy, trustworthiness, and the generalization capabilities of models. [hartmann2023sokmemorizationgeneralpurposelarge] proposed a taxonomy for memorization in LLMs, which was further improved along dimensions of granularity, retrievability, and desirability by [satvaty2025undesirablememorizationlargelanguage]. Based on such classifications, numerous techniques to uncover such behaviour have been presented including methods that utilize injected sequences [huang-etal-2024-demystifying], neuron activations [slonski2024detectingmemorizationlargelanguage], counterfactual reasoning tasks [McCoy2024Embers] or dynamic, prefix-dependent soft prompts [wang-etal-2024-unlocking]. Memorization analyses have also included multimodal behaviour [ju2025watchalbuminadvertentprivacy] and fine-tuned models [zeng-etal-2024-exploring] in recent studies.
Self-knowledge in LLMs: Most studies on self-knowledge in LLMs assess responses in fixed question-answering tasks, equating the self-knowledge problem to binary classifications of answerability [zhang2024largelanguagemodelsgood, ren2024investigatingfactualknowledgeboundary, wen2024perceptionknowledgeboundarylarge]. Recent studies have diversified to understand self-cognition [chen2024selfcognitionlargelanguagemodels], self-recognition using security questions [davidson2024selfrecognitionlanguagemodels], embedding interpretation [speicher2024understandingmemorisationllmsdynamics], and generation-validation consistency of LLMs [li2023benchmarkingimprovinggeneratorvalidatorconsistency] from a self-knowledge perspective. However, most existing methods currently lack feasible alternatives for closed-source models where intermediate stated cannot be analysed [yona-etal-2024-large]. Our research builds on previous findings to establish a link between self-knowledge overconfidence and another persistent challenge of memorization in LLMs, a connection that, to our knowledge, has not been explored before.
7 Conclusion
Overconfidence in capabilities is a severe problem in LLMs, hampering their trustworthiness. In this research, we show how models very likely draw confidence from ‘knowing’ solutions to develop an inflated perception of their reasoning power. We use a novel experimentation approach towards this goal by generating feasible tasks and checking if slight problem perturbations change self-knowledge assessments to analyse if LLM’s tendency to memorize data actually causes overreach in self-knowledge.
Our findings reveal that all models change their feasibility judgments for slightly perturbed tasks in over 45% of cases, demonstrating a systematic overestimation of their capabilities. Such inconsistent self-assessments strongly indicate that LLMs rely on memorized solutions and data to perceive their self-knowledge rather than exhibiting genuine reasoning awareness. Even advanced models struggle to maintain a stable and consistent assessment of their reasoning abilities, particularly in STEM domains, exhibiting high variability even across minimally altered tasks. Memorization-driven overconfidence and instability in self-knowledge are deeply interconnected issues as well, and we urge careful refinement of future training data and model architectures to ensure a balanced self-knowledge stance.
We release our results and evaluation pipeline publicly and hope that researchers can use our method to ensure more trustworthy and dependable LLM-powered applications.
Limitations
  1. 1.
    Catered to STEM domains: Our methodology and prompts are catered for experimentation and evaluation for the STEM domain, where tasks involve structured data. Expanding the pipeline to handle tasks without clear instruction-data separation is a promising area to build on in our research.
  2. 2.
    Small sample size and custom metrics: Since our experiments rely on LLM-generated tasks, which may eventually repeat, we use just over 400 tasks per model to ensure diversity and distinctiveness, although the size may be perceived as relatively small compared to other studies. Similarly, our metrics are defined anew owing to the uniqueness of the exploration. We plan to conduct more exhaustive testing on more models, too, in future work. Also, we plan to establish standard benchmarks and metrics for such evaluation.
  3. 3.
    Multilingual and multimodal experimentation: While we try to incorporate multilingual support in the instruction translation module, we keep our entire prompting and instructing methodology in English only. Expanding our methodology to additional languages and multimodal sources presents a valuable direction for future research.
  4. 4.
    Improved prompt engineering: While our prompts incorporate the majority of good aspects of prompt construction, we do not claim them to be the definitive standard for testing model capabilities. Further refining prompts to better reveal memorization-driven patterns can be an impactful area of improvement.
Appendix
.1 Prompt formats
This section presents the format of all the prompts we use in our experimentation. The prompt format used to generate and validate feasible tasks in STEM domains is shown in Figures 6 and 7, respectively. The prompt used for the ontology replacement module is shown in Figure 8. During task classification, the model is guided to answer only if it deems the task to be feasible; otherwise, it is asked to provide an explanation for infeasibility, as described in the prompt in Figure 9.
.2 Model parameters
We report the generation parameters for all models used in our experiments to ensure transparency and reproducibility. All models were accessed through provider APIs, and common parameter values set during task generation and task classification are listed in Tables 3 and 4, respectively. The model sizes of all closed-source models are approximate and taken from [abacha2025medecbenchmarkmedicalerror].
OpenAI Models: We run our experiments on the flagship models, GPT-4o (~200B parameters). We use the OpenAI Python SDK to access the models via API, specifying seed=1234| and
DeepSeek Models: We use the recently released model: DeepSeek v3 (~671B parameters). The model was accessed using the OpenAI Python SDK by specifying the DeepSeek URL endpoint and authentication details.
Mistral Models: We experiment with Mistral-Large-Instruct-2411 (~123B parameters). The model was accessed using the official API in the Mistral Python SDK.
Anthropic Models: The Claude 3.7 Sonnet model (~175B parameters) was accessed via the official Anthropic Python SDK package.
Figure 6: Prompt format used to generate feasible tasks in any particular STEM domain
Figure 7: Prompt format used to validate feasible tasks and ensure that models are confident of answering self-generated tasks
Figure 8: Prompt format used for the ontology replacement module in the task perturbation pipeline
Figure 9: Prompt format used to classify the perturbed tasks generated by the task perturbation pipeline
Table 3: Parameters used across all models during task generation
Model Parameters
temperature 1.0
top_p 1.0
max_tokens 8096
frequency_penalty 1.0
presence_penalty 1.0
Table 4: Parameters used across all models during task classification
Model Parameters
temperature 0.0
top_p 1.0
max_tokens 8096
frequency_penalty 0.0
presence_penalty 0.0